Find the solution sets for the following:
x^4 + 2x^2 - 15 = 0

Answer by checkley71(8403)

x^4+2x^2-15=0
(x^2+5)(x^2-3)=0
x^2+5=0
x^2=-5
x=sqrt-5
x=+/-2.236i answer.
x^2-3=0
x^2=3
x=sqrt3
x=+/-1.732 answer.